# Identifying Sub-networks in Neural Networks via Functionally Similar Representations

### v1

@@ 3.2 Search Space
3.3 Search Targets
Known Targets
When the search target, denoted as $T$, is a value from a known function, we can directly compare outputs between representations from each layer and known function output $T$. Representations from each layer can be directly compared with the target via a probe. Popular linear probes can be used to assess the similarity between a target and any layer’s representation. For instance, linear regression can be used to model each target $T$ from each representation candidate $Y$, and the residual error is used as the search criteria between $Y$ and $T$. As discussed previously, to deal with the potentially large search space of functions of the target, a more powerful probe (such as a nonlinear MLP function) may have to be used so that it can detect more complex similarities to $T$.
4 Gromov-Wasserstein Distance as a Similarity Measure
Unknown Targets
When the search target is unknown, functionally similar parts by comparing them cannot be identified by comparison to a predefined set of target functions. Instead, we propose to identify the similarities and subnetworks among the representations at each intermediate layer. Each layer, however, posits a representation that potentially has a different distribution, not to mention even different dimensionality depending on the architectures and layers one considers (viz. mlp_in layer and attention layer in transformer blocks). Consequently, representations across layers may be incomparable using standard distance metrics, such as the $\ell_{p}$ norm amongst others.
To overcome these challenging issues, we propose computing distances between representations produced at the same layer for different inputs and match the vertices of this weighted graph – where each dimension of the representation of the inputs are vertices and the distances indicate weights on the edges – with the vertices of a similarly constructed weighted graph from another layer. Essentially, we assume the representations in a layer are samples of the underlying distribution, and we want the best permutation of representation dimensions in one layer that aligns with vertices in another layer, thereby deriving the inter-layer distance. If this inter-layer distance is low, then the two layers we consider to be functionally similar, given the same input data.
Formally, without loss of generality, let ${\bm{Y}}_{1}\in{\mathbb{R}}^{n\times d_{1}}$ and ${\bm{Y}}_{2}\in{\mathbb{R}}^{n\times d_{2}}$ be representations of $n$ examples from two different layers, where the discrete distributions over the representations are $\mu_{1}$ and $\mu_{2}$ respectively, with dimension $d_{1}$ possibly being different from $d_{2}$. Direct distance computation between them is not reasonable. Instead, we seek to compute a coupling or matching $\pi\in\Pi(\mu_{1},\mu_{2})$ between the $n$ examples in each set such that given the pairwise distances $D_{1},D_{2}\in{\mathbb{R}}^{n\times n}$ within representations ${\bm{Y}}_{1}$ and ${\bm{Y}}_{2}$ respectively, the sum of differences between the distances of the matched examples is minimized. Loosely speaking, we aim to find a matching that preserves the pairwise distance as much as possible. In particular, we want to minimize the following:
$\rho({\bm{Y}}_{1},{\bm{Y}}_{2},\mu_{1},\mu_{2},D_{1},D_{2})\triangleq\min\limits_{\pi\in\Pi(\mu_{1},\mu_{2})}$ $\sum_{i,j,k,l}(D_{1}(i,k)-D_{2}(j,l))^{2}\pi_{i,j}\pi_{k,l}$
[…]
4.1 Justification for GW Distance as a Functional Similarity Measure
Theorem 4.1.
The computed GW distance represents the minimal distance over all possible transportation plans between two sets of points from different spaces. In our context, we can also view GW as a measure that quantifies the distance between distance-based (i.g., Euclidean-distance) graphs, with a set of points as its nodes. Hence this would be low if the graph undergoes (nearly) isomorphic transformations between layers. Conversely, a high GW distance indicates a non-distance preserving transformation across layers, potentially reflecting a highly non-linear operation. While GW distance does not reveal the exact function operation, it highlights specific layers for further investigations.
Favorable Properties of GW. Besides the above noteworthy property of GW to map between different spaces, it also has other favorable properties [44, 7]: i) It is symmetric and satisfies triangle inequality. ii) It is invariant under any isometric transformation of the input, which is advantageous because we do not want rotations and reflections to affect our similarity search. This invariance also includes permutation invariance, which is a beneficial property since we want the distance between layer representations to remain unaffected by permutations within the representations in each layer. iii) GW is scalable since it does not require estimating high-dimensional distributions, which is often the case with intermediate representations in large models; instead, it only compare them to obtain a distance measure. iv) GW is monotonic in (positive) scaling of pairwise distances, and hence the same layers should appear to be closer than others even with scaling.
r
(a) Pretrained
(b) Fine-Tuned
(c) tSNE projection
Figure 2: Histogram on pairwise distances on Yelp, for outputs from all transformer blocks in $a)$ a pre-trained BERT model, and $b)$ a fine-tuned BERT model trained on the dataset. $c)$ shows the neighborhood of tSNE projection for a batch of data across layers in the fine-tuned BERT.
Distance Distributions. As an illustrative example, we plot the histogram on pairwise distances for a batch of samples across all transformer blocks in pretrained and fine-tuned BERT models from the YELP review dataset in Figure 2(a) and  2(b). For more details on YELP, we provide a comprehensive discussion in the experiments section  5.2. The results show the distributions on pairwise distances in the fine-tuned model begin to differ from layer 9, consistent with GW distance observed in Figure 6 suggests that significant transformations occur and can be effectively captured by GW.
Neighborhood Change. Complementary to the distribution of pairwise distances, the changing representations of samples could also alter their relative neighborhoods across layers. Using a fine-tuned BERT model, we plot a tSNE projection [38] of representations from a batch of samples on YELP, and visualize it in Figure 2(c). As shown, the sample neighborhood changes across layers, which can be indicative of functional changes that are not captured by comparing distributions alone. However, GW can account for such changes as well. More tSNE figures are shown 9(e) in Appendix D.
Computation Details. We use an existing optimal transport toolbox, pythonot [13], for computing GW distance. Specifically, we use an approximate conditional gradient algorithm proposed in [37], which has a complexity of $O(mn^{2}+m^{2}n)$, where $m$ and $n$ are the dimensions of two spaces (here the number of data samples from two layers being compared). In comparison, the Wasserstein distance may require $O(n^{3}log(n))$ for exact computation, and the Sinkhorn solver [6] operates at nearly $O(n^{2})$. When the dataset is large, we can also sub-sample the dataset to improve the computational efficiency.
5 Empirical Study and Findings
5.1 Synthetic Modular Sum Tasks
Training procedure We train 3 different neural networks with transformer blocks to predict $c$ given $(a,b)$. These networks contain input embeddings for $a$ and $b$, each of size $d$, i.e., $[{\bm{E}}_{a},{\bm{E}}_{b}]\in{\mathbb{R}}^{2d}$, and predict a categorical output $c$ via an unembedding/decoding layer. All parameters in the network are learned. For the first simpler $f_{\text{mod}}$ dataset, we train a neural network consisting of a one-layer ReLU transformer [39], following the same protocol and hyperparameter choices as previous works [30, 45]. We call this Model 0. For the more complex $f_{\text{mod3}}$ dataset, we train two neural networks consisting of three-layer ReLU transformers, with 3 transformer blocks corresponding to the three levels of modular sum functions, and 4 attention heads within each block. The first network, which we call Model E, employs an end-to-end training procedure to directly learn output $c$ given input $(a,b)$. For the second network, which we call Model L, we use the same architecture as Model E but with a layer-wise training approach instead of end-to-end training. Specifically, we use the following procedure:
1. We train the first transformer block layer of Model L to predict $(c_{1},b)$ using an additional linear layer on top, given inputs $(a,b)$.
2. Once block 1 is fully trained, we discard the linear layer, freeze everything before the linear layer, and use its representations of $(c_{1},b)$ to train the second block to predict $(c_{2},b)$, again incorporating an additional layer on top.
3. Finally, we repeat the above step by freezing the first and second block and training the last block to predict $c$, using representations of $(c_{2},b)$.
In all these models, we are able to achieve $100\%$ prediction accuracy. Note that extra linear layers can also be considered as probes but used in training. More details can be found in appendix A. We use the GW measure on Model L with layer-wise training to verify if there is consistency between GW distances and known output $c$’s. To evaluate the capability of handling different dimensions, we directly measure GW distance between 93 intermediate representation $Y$ (see appendix A for search space details) and $c$’s. To speed up computation of GW distance, we randomly sub-sample $1000$ data from a total of $3600$ samples, reducing time from $2$ min to $5$ seconds for each computation.
Table 2: Gromov-Wasserstein Distance Results for Various Targets, for $f_{\text{mod3}}$ dataset.
Model L GW-D for Top Similar Layers $D_{\text{min}}=$
$c_{1}$ Resid-Post1 $0.02$
$c_{2}$ Resid-Post2 $0.03$
$c$ Resid-Post2, Resid-Pre2, Resid-Post3, and 6 others $0.04$
Results The results are shown in the Table 2. We see that in the Model L, the GW distance correctly identifies the most similar layers in accordance with different intermediate $c$’s. The final target $c$ contains 9 similar layers all with distance around $0.04$. In Appendix B, we also test probes since the targets are known. Results shows GW distance can be a reliable alternative to the probes.
Figure 3: Model L (layer-wise training) pairwise GW distance, on $f_{\text{mod3}}$ dataset.
Figure 4: Model E (end-to-end training) pairwise GW distance, , on $f_{\text{mod3}}$ dataset.
Moreover, as previously mentioned GW distance can naturally compare representations across layers with different dimensions. In Fig 4 and Fig 4, we visualize the pairwise GW distance between layer representations without a target for Model L and Model E. Looking at Model L we see predominantly 3 groupings of layers: i) layers roughly from $20$ to $44$ are similar to each other and to layers $52$ to $72$, ii) layers roughly $12$ to $19$ are similar to each other and layers $45$ to $51$ and iii) the initial and last few layers are mainly similar to themselves. Interestingly, the number of groupings corresponds to the 3 functions trained layer-wise in Model L. We also observe differences in patterns across Model L and Model E, suggesting layer-wise and end-to-end training return different networks. Compared to the fixed layer-wise training, end-to-end training in Model E may learn faster in the earlier layers and may not have much to learn in later layers, as the function may not be particularly challenging for it. This could explain why, starting from layer 64, all layers in Model E exhibit similar representations. Moreover, magnitudes of the distances are also different, with Model L showing larger distances, indicating that learning the targets $c_{1},c_{2}$ result in more functional differences. One possible explanation could be that Model E directly operates in the trigonometry space [30], without having to predict the exact integer values until later, thereby suppressing the distances.
Figure 5: Model 0 Pairwise GW distance, , on $f_{\text{mod}}$ dataset.
To gain a deeper understanding of the operations within each transformer block, we visualize pairwise GW distances among layers for Model 0 for dataset $f_{\text{mod}}$ in Figure 5. In this case, we have a total of $31$ representations since only one transformer block is used. We notice the first major difference occurs between layers $13$ and $16$, which are $4$ Attn-Pre (computing key and value product). The second difference occurs between layers $17$ and $20$, which are the first $3$ Attn (computing ${\bm{A}}({\bm{X}})$). This suggests that major computation seems to be done by the attention mechanism. Note that distances are not monotonically increasing across layers, which is expected as the representation spaces can change significantly given the heterogeneity of the operations such as those performed by residual connections and attention within a transformer block.
5.2 NLP Tasks
Setup We now apply GW distance to real natural language processing tasks. We experiment on benchmark sentiment analysis datasets, Yelp reviews 111 https://www.kaggle.com/code/suzanaiacob/sentiment-analysis-of-the-yelp-reviews-data and Stanford Sentiment Treebank-v2 (SST2) from the GLUE NLP benchmark [41], with the goal to predict of the text has positive or negative sentiment, and analyze how different layers from fine-tuning BERT(-base) [8] models perform on these datasets. We use the pretrained BERT to generate 4 fine-tuned models, corresponding to a dense model and 3 sparse models with sparsity levels of $25\%$, $70\%$ and $95\%$ using a state-of-the-art structured pruning algorithm [9]. Training details are in Appendix C. In these BERT models, there are $12$ transformer blocks and we compare the ${\bm{X}}_{\text{in}^{1}}$ and $Y^{l}={\bm{X}}_{\text{out}}^{l}$, for $l=\{1,\dots,12\}$ layers.
Results In the last row of Figure 6(a), we see that the pre-trained BERT does not have major differences among layers, which is not surprising given its accuracy on YELP is only 49.3% (roughly equivalent to random guessing). In the last row of Figures 6(b) to 6(e), we see an interesting pattern emerge, revealing two-to-three major block structures in the (sparse) fine-tuned BERT models identified by our approach. The first major differences occur at block $9$ and then the last three layers ($10$, $11$, $12$) seem to form a distinct block. This seems to indicate that most of the function/task fitting occurs at these later layers. This may be explained by the fact that gradient values are potentially larger at these later layers so they change faster from the pretrained models. The presence of block structures in the GW-distance matrices indicates major functional changes may concentrate at these transition layers. This finding may suggest that for other downstream tasks, we may consider using the representation from layer $8$ and only fine-tuning the later layers. We validate this observation in appendix E.
(a) Pre-trained
(b) Dense
(c) 25% Sparse
(d) 70% Sparse
(e) 95% Sparse
When sparsifying these models, we observe the more sparse models have lesser differences among the layers (with 95% sparse model in 6(e) having the least differences). This is expected as fewer parameters contribute to the final function output besides others being suppressed. Nonetheless, a similar pattern persists, indicating that later layers differ significantly from earlier ones. This observation is consistent with fine-tuning and sparsification literature [23, 9], where it has been observed that later layers typically undergo substantial changes during fine-tuning as they focus on task-specific solutions, while the earlier middle layers remain stable as they capture syntactic and semantic patterns of the language necessary for various tasks. In appendix F, we further investigate the GW distance between layers from different models, providing insights into how representations vary across architectures.
We also compare the proposed GW distance with Euclidean, Cosine, and Wasserstein distance as baselines in Figure 6, on the same YELP dataset and with the same settings. Euclidean distance between two layers’ outputs, shown in the first row of Figure 6, can be seen as the GW distance with a fixed identity-mapping transportation plan for each sample. This validates the low-valued diagonal elements. Off-diagonal elements show greater variation, and it is less obvious there are two distinct sub-groups within layers. The similar pattern is also observed with Cosine and Wasserstein distances, with similar strong diagonal pattern but more pronounced block structures than Euclidean distance. On SST2 dataset, we also observe very similar patterns with the GW distance and 3 baselines, for which we refer the readers to appendix G for detailed results. In both datasets, low distance measure are consistently observed in the diagonal elements, but the overall block structures are not as obvious in the baselines as they are with GW distance, highlighting the effectiveness of the GW distance.
Clustering Besides visualization above, one can also utilize clustering methods to automatically identify the subnetworks from the GW distance. We tested spectral clustering [40] on a similarity matrix computed as the reverse pairwise GW distance matrix. This method successfully identified 2 groups with layers $1\sim 8$ and layers $9\sim 12$.
5.3 Emergence of Subnetworks During Training
We also visualize the GW distance between layers while fine-tuning the pretrained BERT model on YELP datasets in the entire training process, in order to observe when these subnetwork structures begin to emerge. Figure 7 show a few visualization on GW distances at selected training iterations. Layers distances are low in the beginning (observed in Figure 6(a)), but by iteration 300 the last layer begin to differ from other layers. As training progresses, layers 9, 10, and 11 begin to show at iteration 3k and 15k. These growing differences in GW distance reflect the model’s increasing F1 score on the test data. Overall it show the gradual specialization of layers into distinct sub-networks, with each sub-network potentially focusing on different aspects of the task.
(a) Iteration = $300$
(b) Iteration = $3$k
(c) Iteration = $15$k
(d) Iteration = $26$k
We also plot the mean GW distance of all layer pairs in Figure 8. Figure 8a show the mean GW distance over training iterations, and show it grows over time. Figure 8b shows that mean GW distance versus two different accuracy metric on the test dataset. GW distance grow slowly at first, followed by a rapid increase as the model achieves better accuracy and F1 scores. Such observation is consistent with existing “grokking" behavior, where validation accuracy can suddenly increases well after achieving near perfect training accuracy [30]. Similarly, Figure 8c shows a rapid increase in mean GW distance in order to achieve a lower training loss.
6 Discussion
We proposed a novel approach to model interpretation based on functional similarity within intermediate layers of neural networks, using Gromov-Wasserstein (GW) distance to compute such similarities. To the best of our knowledge, our application of GW distance in this context is novel. On algebraic and real NLP tasks, we identified the existence of major sub-components amongst layers, corresponding to functionally meaningful abstractions. Overall, our method provides an automatic low-cost approach to find sub-components within neural networks, facilitating human understanding. Future work could investigate additional models to observe general trends and applications of functional similarity to additional models and tasks. Theoretical study of other properties of GW distances within the context of neural network interpretability is also an interesting future direction.
Limitations and Broader Impact
Acknowledgments and Disclosure of Funding
Appendix A Modular Sum Experiment Details
@@
Appendix B Probes on Modular Sum Dataset: When Target is Known
@@ Results
Appendix C Real NLP Experiment Details
@@
Appendix D Alignment from GW distance
@@
Figure 9: tSNE projection on intermediate representations on Yelp, across BERT models with different sparsity levels. Different Rows: Euclidean, Cosine, Wasserstein, and the proposed GW distances, from top to bottom. Different columns: first column is the pre-trained BERT and the rest are fine tuned BERT models with increasing sparsity (dense, $25\%$, $70\%$ and $95\%$ sparsity).
@@
To complement Figure 2 on other fine-tuned BERT models on YELP, we also plot all the histograms across 5 models in Figure 11.
@@
Appendix E Fine-tuning with Different Layers
@@
Appendix F Cross Model Comparison

### v2

@@ Appendix A Representations in Transformer-based and Convolution Neural Network
We also consider convolution neural networks for computer vision datasets. Specifically, we use a relatively lightweight ResNet9 [17, 40]. The exact locations of the candidate representations considered are listed in Table 3.
Table 3: All representations $Y$ considered in ResNet 9 in experiments.
(Module 0)
Name 0.Conv2d 0.BatchNorm 0.ReLU
Details in-channel = 3, out =64, kernel size = (3,3) Batch Normalization activation
(Module 1)
Name 1.Conv2d 1.BatchNorm 1.ReLU
Details in-channel = 64, out =128, kernel size = (5,5) Batch Normalization activation
(Module 2 & 3: Residual Block )
Name 2.Conv2d 2.BatchNorm 2.ReLU
Name 3.Conv2d 3.BatchNorm 3.ReLU
Details in-channel = 128, out =128, kernel size = (3,3) Batch Normalization activation
(Module 4)
Name 4.Conv2d 4.BatchNorm 4.ReLU 4. MaxPool
Details in-channel = 128, out =256, kernel size = (3,3) Batch Normalization activation Kernel (2,2)
(Module 5 & 6: Residual Block )
Name 5.Conv2d 5.BatchNorm 5.ReLU
Name 6.Conv2d 6.BatchNorm 6.ReLU
Details in-channel = 256, out =256, kernel size = (3,3) Batch Normalization
(Module 7)
Name 7.Conv2d 7.BatchNorm 7.ReLU 7. MaxPool
Details in-channel = 256, out =128, kernel size = (3,3) Batch Normalization activation Adaptive
(Module 8)
Name 8.Linear (classification)
Details in-feature = 128, out =10
Appendix B Modular Sum Experiment Details
@@
Appendix C Probes on Modular Sum Dataset: When Target is Known
@@ Results
Appendix D Baseline Comparison Results on Modular Sum
We have also tested a few baselines that can handle different space dimensions, shown in Figure 8. RSA and CKA reveal different levels of subnetworks within attention layers and across transformer blocks. AGW demonstrates the highest sensitivity to attention computations, while RSM finds the last few layers within each transformer block.
Figure 8: Pairwise (layer) distances on Modular Sum dataset, with layer-wise trained models. Different figures from left to right, top to bottom: RSA, RSM, CKA, MSID, AGW, and the proposed GW distance.
Appendix E Real NLP Experiment Details
@@
Appendix F GW Justification and Alignment
Distance Distributions. As an illustrative example, we plot the histogram on pairwise distances for a batch of samples across all transformer blocks in BERT models from the YELP review dataset in Figure 2. The results in Figure 2 show the distributions on pairwise distances begin to differ from block 9, consistent with GW distance observed in Figure 4, suggesting that significant transformations occur and can be effectively captured by GW.
Neighborhood Change. Complementary to the distribution of pairwise distances, the changing representations of samples could also alter their relative neighborhoods across transformer blocks. We plot a tSNE projection [47] of representations from a batch of samples on YELP, and visualize it in Figure 2b and Figure 9(e). The Jaccard similarity, measuring the overlap between top-5-neighbors of 3 selected samples across different transformer blocks, ranges from 0.0 to 0.43, with average values of $\{0.27,0.26,0.26\}$. The full details are shown in Table 5, as discussed below. Hence, the sample neighborhood changes across blocks, which can be indicative of functional changes that are not captured by comparing distributions alone. However, GW can account for such changes as well.
@@ Appendix F GW Justification and Alignment
Figure 9: tSNE projection on intermediate representations on Yelp, across BERT models with different sparsity levels. Different Rows: Results from all 12 transformer blocks, from top to bottom. Different columns: first column is the pre-trained BERT and the rest are fine tuned BERT models with increasing sparsity (dense, $25\%$, $70\%$ and $95\%$ sparsity).
We also show Jaccard similarity measure on top-5-neighbors, per Euclidean distances on tSNE projection, of each of 3 samples across different transformer blocks. Jaccard similarity is a measure of two sets, computed as their intersection divided by their union. Results are shown in Table 5. This further shows the sample neighborhood changes across layers, and representation similarity measures should account for such changes.
Table 5: Jaccard Similarity on top-5-neighbors of Selected Samples across all transformer blocks.
Sample 1 1 2 3 4 5 6 7 8 9 10 11 12 Mean
Block 0 v.s. 0.25 0.25 0.25 0.11 0.43 0.11 0.25 0.25 0.11 0.11 0.25 0.25 0.27
Sample 2 1 2 3 4 5 6 7 8 9 10 11 12 Mean
Block 0 v.s. 0.11 043 0.11 0.11 0.11 0.0 0.11 0.25 0.25 0.43 0.25 0.25 0.26
Sample 3 1 2 3 4 5 6 7 8 9 10 11 12 Mean
Block 0 v.s. 0.0 0.11 0.43 0.25 0.25 0.25 0.11 0.66 0.11 0.11 0.11 0.0 0.26
@@
To complement Figure 2 on other fine-tuned BERT models on YELP, we also plot all the histograms of pairwise distances between two samples in a batch, across all layers for each of 5 models in Figure 11. Pre-trained models are publicly available models training on other datasets. Row b) to e) are the fine-tuned models on YELP, with different sparsity levels. As one can see, pretrained models do not have much differentiations across layers in the histograms, with maximal KL-divergence of 0.11 between histogram in consecutive layers. Fine tuned models, on the other hard, show larger KL-divergence values, in particular in later layers. For example, Layers 9 in the Dense BERT model contains KL distance of 1.58 from its previous layer. The results show that significant transformations in pairwise distances occur across layers and such distances would be captured by GW distances, as show in Figure 4 and Figure 13.
@@
Appendix G Fine-tuning with Different Layers
@@
Appendix H Baseline Methods and Implementation Details
Besides the standard Euclidean and cosine distances, we compare a few other baselines, as discussed below.
Wasserstein Distance [13]: We use the POT, python optimal transport library pythonot [14], with the algorithm proposed in [2].
Representational similarity metric (RSM) [23]: RSM compares two different spaces by using the $L_{2}$ norms on differences in inter-instances distances. This can be seen as approximation to GW using the fixed and identity transportation plan (i.e., the samples map to itself). We use existing implementation at: https://github.com/mklabunde/llm_repsim/blob/main/llmcomp/measures/rsm_norm_difference.py.
Representational Similarity Analysis (RSA) [23]: RSA is similar to RSM but use correlation instead of $L_{2}$-norm to compute the final distance. Implementation at: https://github.com/mklabunde/llm_repsim/blob/main/llmcomp/measures/rsa.py
Canonical Correlation analysis (CCA) [34]: CCA compute distances based on variances and covariances. Implementation at: https://github.com/google/svcca/blob/master/cca_core.py
Centered Kernel Alignment (CKA) [24]: CKA is based on normalized Hilbert-Schmidt Independence Criterion (HSIC). Implementation at: https://github.com/mklabunde/llm_repsim/blob/main/llmcomp/measures/cka.py
Multi-Scale Intrinsic Distance (MSID) [46]: MSID compute the intrinsic and multiple distance, and can be considered as a lower bound of the GW distance. Implementation at: https://github.com/xgfs/imd/blob/master/msid/msid.py. We have explored different hyperparameter settings with different neighbors k (5 or all batch data available) and number of iterations for SLQ, but results are all similar to the default parameter setting.
Augmented GW (AGW) [7]: AGW considers feature alignment in addition to sample alignment. Its overall objective can be seen as a penalized GW distance. Implementation at: https://github.com/pinardemetci/AGW-AISTATS24/tree/main.
For all methods, we use default parameter settings to obtain results in the paper. Note that RSM, RSA, CCA, MSID, and AGW, along with our proposed approach can handle different dimensions of inputs.
Gromo-Wasserstein Distance [13]: We use the POT, python optimal transport library pythonot [14]. We use the solver based on the conditional gradient [45].
Appendix I More Baselines on YELP
Due to the page limit, here we include baseline results on Yelp Datasets in Figure 12 and Figure 13.
We compare the proposed GW distance with Euclidean, Cosine, and Wasserstein distance as baselines in Figure 12, on the same YELP dataset and with the same settings. Euclidean distance between two layers’ outputs, shown in the first row of Figure 12, can be seen as the GW distance with a fixed identity-mapping transportation plan for each sample. This validates the low-valued diagonal elements. Off-diagonal elements show greater variation, and it is less obvious there are two distinct sub-groups within layers. The similar pattern is also observed with Cosine and Wasserstein distances, with similar strong diagonal pattern but more pronounced block structures than Euclidean distance. we also include 6 other baseline similarity measure in Figure 13. Overall, CKA produces also similar block structures to the proposed GW distance, though with greater variability within block structures. In contrast, other baselines fail to reveal such clear block structures.
(a) Pre-trained
(b) Dense
(c) 25% Sparse
(d) 70% Sparse
(e) 95% Sparse
(a) Pre-trained
(b) Dense
(c) 25% Sparse
(d) 70% Sparse
(e) 95% Sparse
Figure 13: Pairwise (layer) distances on Yelp, across different BERT models. Different Rows: RSA, RSM, CKA, CCA, MSID, AGW, and the proposed GW distance, from top to bottom. Different columns: first column is the pre-trained BERT and the rest are fine tuned BERT models with increasing sparsity (dense, $25\%$, $70\%$ and $95\%$ sparsity). As one can be seen, GW clearly demarcates the (functional) sub-network blocks.
Appendix J Cross Model Comparison